OpenStudy (anonymous):

1. Which is the equation of a circle with center (–2, 3) and radius r = 5? (x + 2)² + (y – 3)² = 10 (x + 2)² + (y – 3)² = 25 (x – 2)² + (y + 3)² = 10 (x – 2)² + (y + 3)² = 25 2. A circle with center (–1, 2) passes through point (2, –2). Which is true? The radius is. The diameter is 10. The equation is (x + 1)² + (y – 2)² = 10. The circumference is 25. 4. Which is the equation of a circle with diameterwith A(5, 4) and B(–1, –4)? (x – 5)² + (y – 4)² = 10 (x + 5)² + (y + 4)² = 100 (x – 2)² + y² = 25 (x + 2)² + y² = 5

OpenStudy (mertsj):

In the equation of a circle: \[(x-h)^2+(y-k)^2=r^2\] What does (h,k) represent? What does r represent?

OpenStudy (mertsj):

Don't you have a book?

OpenStudy (mertsj):

(h,k) is the center. r is the radius. So in your first example, \[(x+2)^2+(y-3)^2=10\] We would rewrite that like this so that it fits the form I posted: \[(x--2)^2+(y-3)^2=(\sqrt{10})^2\] So now we see that h is -2, k is 3 and r is sqrt(10) That means that the center of the circle is (-2,3) and the radius is sqrt(10) What is the center and radius of the second choice?

OpenStudy (anonymous):

Close but not quite.

OpenStudy (anonymous):

is it (x-2)+(y+3)=25

OpenStudy (mertsj):

Use the distance formula to find the radius.

OpenStudy (anonymous):

hey do you have the answers to the other questions?

OpenStudy (anonymous):

Did anyone get these ??

OpenStudy (anonymous):

Heyy! Did anyone get the second or third ont here?? @Danielle96 @Azteck